Two numbers are 20% and 25% lesser than a third number. By how much percent is the second number to be enhanced to make it equal to the firstnumber

Grade:10

1 Answers

Arun
6 years ago
let the 3 rd number be 100
then 25% less than 100 will be ((100-25)/100)*100=75
similarly 20% less will be 80
percent is the second number of the first=(75/80)*100=93.75%